Ask A baseball player hits a home run over the left-field fence, which is 104 m from home plate. The ball is hit at a point 1.09 m directly above home plate, with an initial velocity directed 27.5e above the horizontal. By what distance does the baseball clear the 3.00 m high fence, if it passes over it 3.05 s after being hit?
Answer #1

Solution,

Velocity, v = x/t cos(theta)

V = 104/(3.05 x cos27.5) = 38.44 m/s

After 3.05 sec,

h = ut - 0.5gt^2

h = (38.44 sin27.5) x 3.05 - 0.5 x 9.8 x 3.05^2

h = 8.56 m

Height from ground, 8.56 + 1.09 = 9.64 m

Clearance = 9.64 - 3 = 6.64 m

Comment in case any doubt please rate my answer ....